We selected 3 seeds out of a bag. The bag contains 6 seeds for yellow flowers, 4 red and seven white.
What's the probability that the 3 seeds grow white flowers?
What's the probability that the 3 seeds grow same colored flowers?

Answer by checkley77(12569)

White seeds:
7/17*6/16*5/15=210/4080=7/136 chances (or 5.147%) of the 3 being white seeds.
6/17*5/16*4/15=120/4080=4/136 chances (or 2.941%) of the 3 being yellow.
4/17*3/16*2/15=24/4080=8/1360 chances (or 0.588%) of the 3 being red.
7/136+4/136+8/1360
(70+40+8)/1360
118/1360=.086765 or 8.6765% success in all 3 being the same color.